OpenStudy (anonymous):

Can someone explain to me how the right equation simplifies into the left equation.

OpenStudy (anonymous):

\[-30\pm \sqrt{30^{2}+20(60-h}/-10........3\pm \sqrt{3^{2}+.2(60-h)}\]

OpenStudy (anonymous):

Sorry left equation into the right equation

OpenStudy (aum):

-30 / -10 = 3 For the second part, take -10 into the square root, by squaring -10 which becomes 100. 30^2 / 100 = 30*30/100 = (3^2 * 10^2) / 10^2 20(60-h) / 100 = 0.2(60-h)

OpenStudy (anonymous):

Thank you

OpenStudy (aum):

You are welcome.